# RGE effects on new physics searches via gravitational waves

### v2

@@ Appendix B One-loop improved $\overline{\rm MS}$-parameters
From eqs. (2.20)(2.23), the $\overline{{\rm MS}}$-parameters, e.g., eqs. (2.14) and (2.15), are improved at the one-loop level and are expressed in terms of the physical parameters as follows:
@@ C.1 Higgs, NG, and gauge bosons
±$\pm\sqrt{m_{\chi}^{2}\left(m_{\chi}^{2}-({\overline{g^{\prime}}}^{2}\xi_{B}+\overline{g}^{2}\xi_{W})\phi^{2}-C_{HD}\phi^{2}m^{2}_{\chi}+2C_{HB}\phi^{4}{\overline{g^{\prime}}}^{2}\xi_{B}+2C_{HW}\phi^{4}\overline{g}^{2}\xi_{W}\right)}\bigg{)},$ (C.24)
@@ Appendix D Thermally resummed masses
±$\pm\sqrt{m_{\chi,{\rm res.}}^{2}\left(m_{\chi,{\rm res.}}^{2}-(\bar{g^{\prime}}^{2}\xi_{B}+\bar{g}^{2}\xi_{W})\phi^{2}-C_{HD}\phi^{2}m_{\chi,{\rm res.}}^{2}+2C_{HB}\phi^{4}\bar{g^{\prime}}^{2}\xi_{B}+2C_{HW}\phi^{4}\bar{g}^{2}\xi_{W}\right)}\Bigg{]},$ (D.5)

### v4

@@ Appendix B One-loop improved $\overline{\rm MS}$-parameters
From eqs. (2.20)–(2.23), the $\overline{{\rm MS}}$-parameters, e.g., eqs. (2.14) and (2.15), are improved at the one-loop level and are expressed in terms of the physical parameters as follows:
@@ C.1 Higgs, NG, and gauge bosons
$\pm\sqrt{m_{\chi}^{2}\left(m_{\chi}^{2}-({\overline{g^{\prime}}}^{2}\xi_{B}+\overline{g}^{2}\xi_{W})\phi^{2}-C_{HD}\phi^{2}m^{2}_{\chi}+2C_{HB}\phi^{4}{\overline{g^{\prime}}}^{2}\xi_{B}+2C_{HW}\phi^{4}\overline{g}^{2}\xi_{W}\right)}\bigg{)},$ (C.24)
@@ Appendix D Thermally resummed masses
$\pm\sqrt{m_{\chi,{\rm res.}}^{2}\left(m_{\chi,{\rm res.}}^{2}-(\bar{g^{\prime}}^{2}\xi_{B}+\bar{g}^{2}\xi_{W})\phi^{2}-C_{HD}\phi^{2}m_{\chi,{\rm res.}}^{2}+2C_{HB}\phi^{4}\bar{g^{\prime}}^{2}\xi_{B}+2C_{HW}\phi^{4}\bar{g}^{2}\xi_{W}\right)}\Bigg{]},$ (D.5)